### v1

# TuRTLe: A Unified Evaluation of LLMs for RTL Generation

Abstract
The rapid advancements in LLMs have driven the adoption of generative AI in various domains, including Electronic Design Automation (EDA). Unlike traditional software development, EDA presents unique challenges, as generated RTL code must not only be syntactically correct and functionally accurate but also synthesizable by hardware generators while meeting performance, power, and area constraints. These additional requirements introduce complexities that existing code-generation benchmarks often fail to capture, limiting their effectiveness in evaluating LLMs for RTL generation. To address this gap, we propose TuRTLe, a unified evaluation framework designed to systematically assess LLMs across key RTL generation tasks. TuRTLe integrates multiple existing benchmarks and automates the evaluation process, enabling a comprehensive assessment of LLM performance in syntax correctness, functional correctness, synthesis, PPA optimization, and exact line completion. Using this framework, we benchmark a diverse set of open LLMs and analyze their strengths and weaknesses in EDA-specific tasks. Our results show that reasoning-based models, such as DeepSeek R1, consistently outperform others across multiple evaluation criteria, but at the cost of increased computational overhead and inference latency. Additionally, base models are better suited in module completion tasks, while instruct-tuned models perform better in specification-to-RTL tasks.
I Introduction
The rapid advancements in large language models (LLMs), known for their ability to process and generate human-like text, have unlocked new possibilities across a wide range of domains [1, 2]. Domain-specific LLMs have gained significant attention due to their strong performance in specialized tasks, including financial engineering [3], biomedical research [4], and scientific computing [5, 6]. LLMs also assist developers by suggesting code snippets, solving common coding challenges, and providing clear explanations of complex concepts [7, 8, 9].
In the field of Electronic Design Automation (EDA), researchers are increasingly exploring the use of LLMs to accelerate hardware design[10, 11, 12]. In the traditional digital system design flow, engineers invest significant effort in implementing precise functionality using hardware description languages (HDLs). LLM-based solutions aim to bridge this gap by translating functional specifications directly into HDL code, such as Verilog. This approach has the potential to revolutionize hardware design and verification by streamlining Verilog coding, optimizing circuit implementations, and automating time-consuming design tasks [13]. Recent research has explored various techniques for LLM-driven Verilog code generation, including: prompt engineering to enhance model responses [14], training and fine-tuning on Verilog-specific datasets [15, 16, 17, 18, 19], and agent-based approaches to improve iterative refinement and debugging [20, 21, 22, 23, 24]. However, several research questions about the quality and fidelity of the generated Verilog code and, hence, the hardware design, remain largely unanswered. As for code generation for other programming languages, LLM-based solutions need to generate code that is syntactically and functionally correct, i.e., , the generated code must compile and generate the correct results. High-quality code must also provide reasonable performance. EDA tools pose the additional constraints that hardware generators need to be able to take the generated RTL and be able to synthesize it to produce correct hardware components. Additionally, non-functional constraints, such as area and power, play an important role when resources are limited. It is evident that evaluation methodologies and benchmarks designed to assess LLM that produce traditional code do not fully cover the entire evaluation space and are not completely suitable to evaluate LLM-based EDA.
To mitigate this issue and assess the effectiveness of current LLM for EDA, several benchmarks have been introduced, including VerilogEval[25, 26], RTL-Repo[27], and RTLLM[28]. While these benchmarks evaluate different aspects of Verilog code generation, including single-line completion, single-module generation, and specification-to-RTL translation, none fully evaluates the entire hardware design flow, and some use different metrics. This makes it challenging to compare results across benchmarks and LLMs. A standardized evaluation methodology and infrastructure would allow researchers to systematically assess LLM capabilities, identify strengths and limitations in existing benchmarks and models, and drive further advancements in LLM-driven hardware design automation.
This work introduces TuRTLe, a comprehensive evaluation framework for RTL generation. It integrates multiple benchmarks into a single fully-automated evaluation framework, providing a standardized and comprehensive assessment of code. Coverage is provided for different RTL design goals, including syntax and correctness of the generated code, synthesizability of the hardware circuits produced by hardware generators, together with their corresponding performance, area, and power metrics. These goals are tested under different conditions, in particular single-line completion, module completion, and generation from specifications expressed in natural language. TuRTLe is used to benchmark 13 models, providing a comprehensive, wide and updated review on the state of open models. This includes general purpose models, coding, and RTL-specific LLMs.
Our analysis show that temperature, context length, model structure (base vs. instruct), and reasoning mechanisms all play a crucial role in LLM performance across EDA-related tasks. Base models demonstrate stronger performance in module completion, effectively filling in partial RTL definitions, while instruct-tuned models excel at generating RTL from natural language specifications, leveraging their training in instruction-following tasks. Models employing reasoning prompting chains, such as DeepSeek R1, exhibit remarkable efficiency in handling complex tasks like specification-to-RTL, though at the cost of increased inference time and token generation. Overall, our findings indicate that while current LLMs are highly proficient in generating syntactically correct RTL code, they struggle with functional correctness. However, once functional correctness is achieved, synthesizability is generally not an issue, and the resulting circuit designs are often comparable to human-defined modules.
II Evaluation Methodology
Evaluating LLM-generated RTL code requires a comprehensive and systematic approach that considers the many aspects of the hardware design process. TuRTLe integrates existing benchmarks including VeriGen [15], VerilogEval [25, 26], RTL-Repo [27], and RTLLM [28], within a single infrastructure, allowing for direct comparison of LLM performance across various RTL generation tasks. Focusing on versatility and automation in the construction and running of TuRTLe, we choose a widely adopted framework in the field of code generation (BigCode Evaluation Harness [29]) as a foundation. This tool can easily load a variety of generative models, and already implements multiple LLM tasks and metrics. Focusing on scalability, we fork from the vLLM Code Harness project [^1] https://github.com/iNeil77/vllm-code-harness, a particularly efficient derivative, which facilitates running larger models through the vLLM [30] library. Through these choices, TuRTLe is compatible by design with most popular practices in the field.
This section first reviews three tasks involved in RTL generation that TuRTLe focuses on: single-line completion, module completion, and specification-to-RTL conversion, all in §II-A. Next, it outlines the five RTL design goals assessed by TuRTLe: line-level contextual accuracy, syntax correctness, functional correctness, synthesizability, and post-synthesis quality (see §II-B). To evaluate LLM performance on these goals, we employ a set of metrics, including exact matching, Pass@1, and a novel PPA-Score (detailed in §II-C). Table I shows the relationships between the various tasks, design goals, metrics, and tools used in the TuRTLe framework.
II-A Generation Tasks
Our evaluation framework assesses LLMs across three fundamental RTL code generation tasks, each representing a distinct levels of complexity and scale. The smallest of tasks, in the sense of generation length, is Single-Line Completion (SLC), which focuses on the model’s ability to predict the next line of code given a partial context. This task closely resembles auto-completion scenarios, where engineers rely on coding assistants to streamline development, and evaluates the capacity of models to produce contextually coherent code.
The second task in complexity is Module Completion (MC), which requires LLMs to generate a complete module based on a given function description or module signature. This task evaluates an LLM’s ability to process the behavioral description of a hardware module and produce correct implementations.
Finally, the third and most complex task is Specification-to-RTL (S2R) generation, where the LLM must generate a complete implementation from a natural language hardware specification. This task does not provide a predefined module interface, so models must infer additional information from the specification, such as the module name, the correct ports and parameter types and names. This mimics the process of translating human-readable specifications into working hardware descriptions. Given their fundamental differences in complexity and nature, results for these three approaches are reported separately.
II-B Design Goals
TuRTLe defines five key design goals to evaluate LLM-generated RTL code, ensuring a comprehensive assessment of validity, correctness, feasibility, and efficiency within the hardware design workflow. These goals are aligned with the supported tasks to maintain compatibility. Given SLC, MC and S2R, we consider the following design goals.
Line-level Contextual Accuracy (LCA) tests the ability of the models to produce line completions that are as close as possible to a given reference answer. This goal aligns with the SLC task, and evaluates whether the predicted line is both syntactically correct and contextually coherent within the surrounding logic.
The four remaining goals can all be applied to both MC and S2R tasks. Syntax Correctness (STX) ensures that the generated HDL code complies with the language’s grammar rules and can be processed by standard HDL analyzers, synthesizers, and simulators [31]. The correctness of generated designs is verified by parsing and elaboration, using tools that check for syntax errors, missing constructs, and structural inconsistencies. Functional Correctness (FNC) ensures that the generated HDL code exhibits the expected behavior as specified in the design requirements (prompt) [32, 33]. FNC is verified through behavioral simulation, where the generated design is executed alongside a predefined testbench in an HDL simulator. While this approach provides a practical verification method, it relies on the assumption that testbenches accurately capture functional discrepancies.
Synthesizability (SYN) assesses whether the generated HDL code can be successfully synthesized into a gate-level netlist using a synthesis tool. In practical hardware design, synthesis tools support only a subset of HDL constructs, making synthesizability a key criterion for real-world applicability. SYN is verified by analyzing and synthesizing the generated RTL code using a synthesis toolchain, ensuring that the design is implementable in hardware. This is critical for models that generate RTL intended for manufacturable chip designs.
Finally, Post-Synthesis Quality (PSQ) evaluates the implementation efficiency of the synthesized design based on Power, Performance, and Area (PPA). PPA metrics are widely used in chip design optimization to compare different implementations and ensure that a design meets key hardware constraints [34]. In this work, we use PPA to compare LLM-generated designs with manually optimized reference implementations. We extract area and power directly from the PPA report and evaluate performance based on maximum delay, defined as the difference between the clock period and the worst slack reported by static timing analysis. In this way, all three PPA metrics are represented as positive numbers, where 00 represents the minimum possible value, leaving the maximum unbounded.
TABLE I: Design goals and metrics for benchmarking RTL generation. The use of open-source tools ensures broader accessibility and reproducibility.
Tasks Goals Metrics Tools
SLC LCA Pass@$1$ Exact Match
MC STX Icarus Verilog
S2R FNC
SYN OpenLANE (Yosis)
PSQ PPA-Score OpenLANE (OpenROAD)
II-C Measures and Metrics
To evaluate the five goals described above (LCA, STX, FNC, SYN and PSQ) we use different tools and evaluation measures tailored to each case. For LCA, prior work [27] has explored both exact and fuzzy matching techniques. Exact Matching (EM) determines whether the generated line precisely matches the reference HDL line in the dataset, while Fuzzy Matching relies on edit distance and semantic similarity metrics. Our experimentation shows both metrics are strongly correlated, which is why we report the more precise EM alone.
For MC and S2R tasks, we set up an evaluation pipeline that sequentially tests STX, FNC, SYN, and PSQ. STX is evaluated by compiling a design along with its testbench using Icarus Verilog [35] and checking for errors. If no errors are issued at compile time, FNC is evaluated by running the simulation executable generated by the compiler and checking if the testbench passes. Functionally correct codes are tested for SYN, by elaborating the design with OpenLANE (Yosys) [36]. For the four previous goals (LCA, STX, FNC, and SYN), we report the same metric Pass@$1$. This is a Pass@$k$ metric [37, 38], which measures the probability that at least one of $k$ generated solutions passes the corresponding test criteria. We set $k=1$ to focus on the requirement of generating the correct result on the first try.
For those designs that pass SYN, PSQ is computed by using OpenLANE [39] to synthesize the code into a netlist and extract post-synthesis PPA metrics. Designs are synthesized using the SKY130A open-source PDK [40] with a $10ns$ delay constraint. In this work, we introduce a novel metric called PPA-score to measure PSQ, focusing on the comparison between LLM-generated synthesizable code and a golden, human-crafted reference.
PPA-Score
Given a benchmark with $n$ problems, for each problem $i\in\{1,\dots,n\}$, we generate $m$ candidate solutions. Each generation $j\in\{1,\dots,m\}$ is processed through the evaluation pipeline sequentially: STX, FNC, SYN, PSQ. Each stage only processes results that passed the previous one, creating a cascade score where $STX\geq FNC\geq SYN\geq PSQ$ (failures in previous stages are reported as automatic fails in the next ones). STX, FNC and SYN are binary evaluations, pass or fail, which are aggregated using Pass@$k$. However, PSQ requires not only to aggregate numeric values that have to be analyzed in comparison with the reference PPA of the golden solution, but also to take into account that models that produce more synthetizable code can be evaluated on a larger set of problems, thus increasing the challenge and confidence in the results obtained. To achieve this, let $p_{i,j}$ represent the PPA metric (power, performance, or area) from the LLM for candidate $j$ of problem $i$:
  1. 1.
    Each $p_{i,j}$ is compared against the corresponding PPA value $g_{i}$ of the golden solution. For that, instead of aggregating $p_{i,j}$ we compute $\hat{p}_{i,j}=p_{i,j}/g_{i}\in(0,+\infty)$.
  2. 2.
    For generations that do not pass STX, FNC and SYN evaluations, $p_{i,j}$ cannot be computed. As a result, we set a failure value of $\hat{p}_{i,j}=2$ (e.g. producing a design two times bigger than the human reference in the case of the area metric). This approach also requires us to limit results that pass the previous evaluations but perform worse than this threshold (i.e. $\hat{p}_{i,j}>2$).
  3. 3.
    We divide by 2 to get a range from 0 to 1.
  4. 4.
    We flip the result so that the metric behaves as the rest of the goals (higher is better).
Following these steps, if a design has score = 0 means it requires twice the area, the power or the performance, when compared to the human reference. Any measurement worse than double (i.e., on the negative side of the score) is clipped to zero (because we set $\hat{p}_{i,j}=2$). A score = 0.5 are designs with an area, power or performance equal to that of the human reference. Finally, a score = 1 can only be obtained by chips which occupy no space, execute in no time, and consume no energy. The final formula considering all generations of an LLM for a given benchmark is then computed as:
[…]
TABLE II: Design goals covered by RTL benchmarks. Number of designs reports two sizes: Samples for LCA (left) and samples for the rest of goals (right).
Benchmark LCA STX FNC SYN PSQ Num. Designs
RTL-Repo 1,174
VeriGen 17
VerilogEval 156
RTLLM 50
TuRTLe 1,174 223
II-D Integrated Benchmarks
TuRTLe initially integrates a selection of four benchmarks, selected by quality, variety and size. These target specific tasks and align with different design goals introduced in our work. Table II provides a comparative overview of the design goals covered by each integrated benchmarks, together with the number of problem descriptions included. This number is reported separately for LCA (left) and the rest of goals (STX, FNC, SYN and PSQ, right), as they have different requirements.
RTL-Repo [27]
This benchmark is designed to evaluate LLMs’ capabilities in the single-line completion task by assessing their ability to perform local edits within large-scale Verilog projects. It consists of 4,098 Verilog code samples sourced from public GitHub repositories. To construct prompts, multiple Verilog files from a given project are concatenated into a single input, which is then truncated at a predefined context length. The truncated prompt is fed to the LLM, requiring it to predict the next line of code immediately following the truncation point. Moreover, TuRTLe supports both the full dataset (4,098 samples) and the test split (1,174 samples) for benchmarking, but we adopt the test split as the default configuration. The test partition is large enough to produce stable results with minimal variance, reducing the risk of overfitting. More importantly, using only the test set mitigates data contamination risks, enhancing the credibility of the evaluation metric. RTL-Repo primarily targets the LCA goal, which is the evaluation aspect implemented in TuRTLe.
VeriGen [15]
This benchmark evaluates LLMs’ performance in the MC task using 17 Verilog problems categorized into basic, intermediate, and advanced difficulty levels. Basic problems include simple components such as wires and logic gates, while advanced tasks involve more complex designs like finite state machines. For each problem, three levels of prompt specificity are provided: low, medium, and high with each offering different amounts of detail. The low-detail prompts include only the module header, which is generally insufficient for meaningful completion. In contrast, the high-detail prompts provide excessive information about the problem, making them impractical for real-world applications. To ensure a balanced and realistic evaluation, TuRTLe adopts the medium-detail prompts as the standard configuration. VeriGen was previously evaluated for STX, FNC, and SYN goals. In TuRTLe, we extend its scope by incorporating an additional design goal (PSQ) to assess the efficiency of LLM-generated RTL in terms of PPA.
VerilogEval [25, 26]
This benchmark consists of 156 problems sourced from HDLBits [^2] https://hdlbits.01xz.net/wiki/Problem_sets, designed to assess LLMs’ performance in two tasks: MDC and S2R. In the MDC task, models are provided with a problem statement along with a module header and are required to generate the missing body of the module. In the S2R task, models receive a prompt resembling a high-level design specification and must generate an entire module from scratch. VerilogEval was originally designed to evaluate STX, FNC and SYN design goals. TuRTLe extends its evaluation scope by incorporating PSQ. We manually corrected the reference implementation of six problems, to address synthesizability issues as explained in Appendix IV.
RTLLM [28]
This benchmark introduces 50 human-crafted designs of varying complexity, designed to evaluate LLMs on S2R. Originally, RTLLM was developed to assess LLM performance across three evaluation goals: syntax correctness (our SYN), functional correctness (our FNC), and design quality including power, performance, and area metrics (our PSQ). Notice RTLLM definition of syntax correctness is differently from the STX goal used in this work. Specifically, RTLLM evaluates syntax correctness based on whether a design can be successfully synthesized into a netlist without syntax errors which is effectively combining aspects of both STX and SYN as defined in this work. RTLLM originally employs a “success rate metric,” which is conceptually similar to the Pass@$k$ metric [38] used in this work, but it does not strictly adhere to the requirement that $N>k$. We corrected the reference implementation of four problems and excluded two to address functional and synthesizability issues as explained in Appendix IV.
III Experiments
This section presents a comprehensive evaluation of LLMs for RTL generation, analyzing their performance across multiple tasks and design goals using the TuRTLe framework. The following subsections detail the experimental setup, benchmarking methodology, and key findings from our evaluation.
III-A Experimental Setup
TuRTLe evaluates a diverse set of open LLMs, categorized into three groups based on their specialization. The first category comprises general-purpose LLMs, such as LLaMA [41] and DeepSeek [42], which are designed for broad language understanding but lack specific optimizations for code generation. The second category includes coding-specific LLMs, such as QwenCoder [43] and OpenCoder [44], which are fine-tuned on programming tasks. Finally, the third category consists of RTL-specific LLMs, such as CodeV [45] and HaVen [46], which are explicitly optimized to generate Verilog and other RTL constructs more accurately. By evaluating models from all three categories, this study provides a comprehensive understanding of the ability of different type of LLMs for specific RTL generation tasks.
We use the Pass@$1$ metric to measure the probability that a model generates a correct solution on the first attempt. However, its accuracy is dependent on the number of samples ($N$) used per evaluation. To determine the optimal value of $N$, we performed ten independent runs on five selected models using varying sample sizes of $N={1,3,5,10,20}$. As shown in Figure 1, the variance decreases as $N$ increases, confirming that small sample sizes lead to unstable results. While larger $N$ values further reduce variance, they impose significant computational overhead. A sample size of $N=5$ is chosen as the optimal configuration, balancing computational efficiency with statistical stability.
Figure 1: Pass@$1$ variance among ten runs while increasing sample size $N$.
TABLE III: Pass@1 performance (higher is better) with n=5 for line and module completion (top table) and specification-to-RTL (bottom table). Each table is split vertically in three, including general purpose LLMs (top), coding LLMs (middle) and RTL-specific LLMs (bottom).
Line Syntax Functionality Synthesis Post-Synthesis Quality
Models RTL-Repo VerilogEval (MC) VeriGen VerilogEval (MC) VeriGen VerilogEval (MC) VeriGen VerilogEval (MC) VeriGen
Power Perf. Area Power Perf. Area
DeepSeek R1 33.02 97.95 94.12 80.26 60.00 79.62 54.12 39.35 38.16 39.21 26.62 26.97 28.01
Llama 3.1 405B FP8 33.29 91.41 72.94 44.74 45.88 44.10 44.71 21.98 20.74 21.66 19.24 22.19 21.08
Qwen2.5 72B 37.19 81.67 70.59 53.08 27.06 52.56 27.06 26.08 24.92 25.83 12.74 13.50 13.89
StarChat2 15B v0.1 13.24 81.54 92.94 39.36 50.59 38.59 50.59 19.21 18.28 19.05 24.02 25.23 25.73
[…]
QwenCoder 2.5 32B 30.44 84.87 72.94 45.51 41.18 44.87 41.18 22.26 21.48 22.20 20.56 20.67 20.87
DeepSeek Coder 33B 30.58 78.72 83.53 39.49 29.41 38.33 29.41 18.92 18.20 18.76 14.52 14.74 14.67
OpenCoder 8B 16.63 79.87 92.94 36.03 43.53 35.51 37.65 17.57 16.74 17.52 17.19 18.76 19.06
[…]
OriGen DeepSeek Coder 19.45 79.36 87.06 43.08 35.29 42.95 35.29 21.50 20.13 21.33 16.55 17.70 18.35
HaVen-CodeQwen 25.38 93.33 97.65 50.00 48.24 48.72 42.35 23.37 23.39 23.09 20.21 21.15 21.25
CodeV-CL-7B 12.39 91.92 98.82 36.79 44.71 36.41 38.82 18.15 16.88 18.05 19.06 19.38 19.35
CodeV-QW-7B 20.56 93.85 57.65 52.56 25.88 51.15 20.00 25.64 24.22 25.56 9.39 9.99 9.94
Syntax Functionality Synthesis Post-Synthesis Quality
Models VerilogEval (S2R) RTLLM v2.0 VerilogEval (S2R) RTLLM v2.0 VerilogEval (S2R) RTLLM v2.0 VerilogEval (S2R) RTLLM v2.0
Power Perf. Area Power Perf. Area
DeepSeek R1 96.54 91.43 79.74 67.76 78.97 63.27 38.94 37.82 38.76 35.64 31.95 34.50
Llama 3.1 405B Instr. 89.10 65.71 57.05 37.55 56.67 35.92 27.18 27.0 26.79 19.70 16.04 18.91
Qwen2.5 72B 81.15 82.04 51.15 47.35 50.38 46.53 25.40 23.83 24.52 24.83 23.88 25.46
StarChat2 15B v0.1 86.54 85.71 38.72 42.45 38.59 42.45 19.18 17.99 19.00 22.44 21.03 22.53
[…]
QwenCoder 2.5 32B 87.69 79.59 45.64 43.27 43.33 42.04 21.51 20.72 21.17 22.02 20.95 22.03
DeepSeek Coder 33B 57.82 83.67 19.87 43.67 19.87 42.86 9.94 9.83 9.47 23.28 21.19 23.20
OpenCoder 8B 75.77 75.10 28.59 46.53 28.21 42.86 13.81 13.16 13.71 22.24 21.47 21.73
[…]
OriGen DeepSeek Coder 90.26 23.67 46.54 12.65 46.92 10.61 23.38 22.18 23.44 5.33 4.61 4.79
HaVen-CodeQwen 90.26 82.45 45.90 40.41 44.36 38.37 21.77 21.23 21.46 19.10 18.31 18.92
CodeV-CL-7B 55.38 69.80 27.05 37.14 26.79 35.10 13.20 12.39 13.03 18.92 16.88 17.89
CodeV-QW-7B 41.79 71.02 19.10 35.51 18.72 27.76 9.36 9.36 9.38 14.85 12.21 13.78
Figure 2: Average STX (left) and FNC (right) scores on VerilogEval and RTLLM across three model categories (General, Coding, and RTL-Specific) at different temperature settings (0.2, 0.5, 0.8). Darker shades indicate higher scores.
III-B Benchmarking Insights
Temperature Ablation
We examine the impact of stochasticity on LLM performance in RTL generation. Temperature controls the randomness of model-generated outputs, influencing both creativity and precision. To assess its effect, we evaluate model performance across three temperature settings: 0.2, 0.5, and 0.8 for Pass@$1$, focusing on STX and FNC goals. Figure 2 shows temperature = 0.2 yields the highest accuracy across all model categories. This suggests that minimal randomness is optimal for RTL generation tasks. Consequently, 0.2 is set as the default temperature for all subsequent experiments to ensure consistency and reproducibility.
Impact of the Context in the Prompt
We analyze how the length of the context from the Verilog project that is given in the prompt affects performance in the single line completion task. We do this by evaluating six models, spanning general-purpose, coding and RTL-specific, using context sizes of 2,048, 4,096, and 8,192 tokens. Our results show that longer contexts consistently improve performance, confirming that LLMs benefit from additional context length when making predictions. Increasing the context length from 2,048 to 4,096 tokens results in a notable improvement of +3.35 in Exact Match scores with a standard deviation of 1.02. Extending the context further to 8,192 tokens provides an additional boost of +2.40 (standard deviation of 1.16). Based on these findings, 8,192 tokens is selected as the standard context length for the SLC task while evaluation RTL-Repo.
Base vs. Instruct-Tuned Models
Figure 3: Comparison of Base and Instruct-Tuned model performance on VerilogEval for MC and S2R tasks across five model families.
VerilogEval supports both MC and S2R tasks using the same underlying dataset, allowing for a direct comparison of how problem definition and presentation impact model performance. Figure 3 compares the performance of base and instruct-tuned variants across ten models spanning five distinct model families: OpenCoder [44], Qwen 2.5 [47], QwenCoder 2.5 [43], DeepSeek Coder V1 [42], and Llama 3.1 [48]. Our results show that there is a clear trend: base models tend to perform better in MC task, where they leverage learned syntax patterns to complete partial module definitions, whereas instruct-tuned models excel in S2R tasks, benefiting from their training on instruction-following tasks that improve their ability to interpret high-level specifications. Given computational efficiency and space constraints, we report base model performance on MC and SLC task, while instruct-tuned models are evaluated on S2R tasks, ensuring that each variant is assessed in the most relevant scenario for its capabilities.
III-C Model Results
This section presents a detailed evaluation of LLM performance across the five key design goals. Table III-A provides a structured comparison, with the top section reports the performance of various models on LCA and MC tasks, while the bottom section highlights results for the S2R task.
Model Performance across RTL Tasks
DeepSeek R1 emerges as the best-performing model across most evaluation metrics, largely due to its autoregressive reasoning chain, which enables iterative refinement of generated code. This advantage is particularly evident in tasks requiring step-by-step reasoning, such as MC and S2R, where DeepSeek R1 significantly outperforms other open models. However, this improvement comes at the cost of increased computational overhead during inference. On RTLLM, for instance, DeepSeek R1 required an average of 5,606 additional tokens solely for its reasoning chain, leading to substantially longer generations. To accommodate these demands, DeepSeek R1 was evaluated using a context length of 16,384 tokens, whereas other models performed effectively with 2,048 tokens. On the other hand, in scenarios where reasoning chains are not employed, such as in real-time SLC task, DeepSeek R1 performs comparably to other top-performing models, suggesting that its primary advantage lies in multi-step reasoning rather than immediate completion accuracy.
A clear distinction is observed in model specialization. General-purpose models perform well in SLC and S2R tasks, where broad contextual understanding plays a crucial role. On the other hand, RTL-specific models excel in MC tasks, leveraging their domain-specific training to generate functionally accurate Verilog modules. Coding-specific models, such as QwenCoder 2.5 32B, demonstrate stable performance across multiple tasks, balancing generalization with code-optimization capabilities.
Figure 4: Average model performance across design goals for specific-to-RTL task.
Model Performance by Design Goals
Figure 4 shows average model performance across four design goals (STX, FNC, SYN, and PSQ) for specific-to-RTL task evaluated across all benchmarks. A clear performance degradation is observed as the evaluation criteria become progressively challenging. On average, models achieve a relatively high STX score of 73.88%, indicating that they generate syntactically correct Verilog in most cases. However, transitioning from syntax correctness to functional correctness results in a substantial drop, with the FNC score averaging 40.11% which is a sharp decrease of 33.76%. This decline suggests that while models are proficient at producing Verilog code that adheres to syntax rules, they struggle to generate functionally valid implementations. The significant gap between STX and FNC highlights a key limitation of current LLMs.
Regarding PPA, the PPA-score presented in §II-C has the characteristic that each model has a different human baseline, equal to their SYN score divided by 2. This is because all designs that do not pass synthesis are considered as failures, which ensures the consistency of the whole framework (every step is equal or lower than its predecessor), and illustrates the current limitations of LLMs and benchmarks assessing them. Having this into account, it can be seen in Table III-A that all models perform very close to their human baseline. This could be caused by the simplicity in the synthesized designs, as there is limited room for either significant optimization or flagrant mistakes in simple examples. However, this view is still useful as a sanity check, ensuring that the code produced is at least as optimal as the human reference.
Figure 5: Spearman correlation among benchmarks for SYX and FNC goals.
Benchmark Correlation Analysis
Figure 5 presents the Spearman correlation between benchmarks for STX and FNC, which measures the monotonic relationship between rankings, helping to determine how similarly models perform across benchmarks. The highest correlation observed is between VerilogEval S2R and VeriGen, a moderate positive correlation (0.49). The rest of correlations range between that value and the lack of correlation (zero). These results suggest that each benchmark captures different aspects of RTL generation challenges, emphasizing the importance of combining them for obtaining a comprehensive evaluation of LLM-generated Verilog code.
IV Conclusions
This work presents a comprehensive evaluation of open LLMs for RTL generation, assessing their capabilities across various tasks (single-line completion, module completion, and specification-to-RTL) and design goals (line-level contextual accuracy, syntax correctness, functional correctness, synthesizability, and post-synthesis quality). To achieve this, we developed TuRTLe, a unified evaluation framework that integrates multiple benchmarks into a single, automated infrastructure. This framework simplifies experimentation, enables systematic model evaluation, and allows for future extensibility, making it adaptable to new benchmarks and models.
Our results provide key insights into LLM performance for RTL generation. Reasoning-based models, like DeepSeek R1, generally outperform others but incur higher token generation and inference latency. Model structure also influences task performance, as base models excel in module completion, while instruct-tuned models are more effective in specification-to-RTL tasks, benefiting from structured prompt training. Despite achieving high STX, LLMs struggle with FNC design goals, as evidenced by a significant performance drop between these metrics. Many models generate compilable but functionally incorrect Verilog, highlighting a key limitation. To further assess synthesized designs, we introduce a new PSQ metric, which evaluates performance, power, and area, enabling a direct comparison between LLM-generated and human-designed RTL implementations.
Overall, this analysis represents a significant step toward a comprehensive evaluation of LLMs for EDA workflows, providing insights into their capabilities and limitations in RTL generation. However, our findings also underscore the need for more sophisticated and realistic benchmarks that better reflect real-world design challenges, including multi-module architectures and complex module interconnections. The TuRTLe framework source code is released on GitHub [^3] https://github.com/HPAI-BSC/TuRTLe, ensuring reproducibility. A complete leaderboard, including more models and metrics, is hosted in HuggingFace [^4] https://huggingface.co/spaces/HPAI-BSC/TuRTLe-Leaderboard. Moving forward, we plan to expand our framework with additional benchmarks, further refining the evaluation of LLMs in hardware design automation.
Acknowledgment
This work is supported by the AI4S fellowships awarded to Gokcen Kestor, Emanuele Parisi, Razine Moundir Ghorab, Cristian Gutierrez and Miquel Albertí Binimelis as part of the “Generación D” initiative, Red.es [^5] https://www.red.es/es, Ministerio para la Transformación Digital y de la Función Pública, for talent attraction (C005/24-ED CV1). Funded by the European Union NextGenerationEU funds, through PRTR. Additionally, this work has been partially funded by the Generalitat de Catalunya (contracts 2021-SGR-00763 and 2021-SGR-01187), and by the project PID2023-146511NB-I00 funded by the Spanish Ministry of Science, Innovation and Universities MCIU /AEI /10.13039/501100011033 and EU ERDF. We are grateful to the Operations department at BSC for their technical support.